OpenStudy (anonymous):

Which of the following is the formula for the geometric sequence? 320, 80, 20, 5,... A. an = 320(1/4)n-1 B. an = 320(4)n-1 C. an = 4(320)n-1

jimthompson5910 (jim_thompson5910):

First find the common ratio: r = (any term but the first)/(previous term) r = (20)/(80) r = 1/4 So the common ratio is 1/4. Basically, we divide each term by 4 to get the next term The first term is 320. This is a[1]. So a[1] = 320 In general, the geometric sequence has the following form a[n] = a[1]*r^(n-1) Plug in the values we found above to get a[n] = 320*(1/4)^(n-1) So the answer is choice A